40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 39 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  15036   Tue Nov 19 21:53:57 2019 gautamUpdatePEMFollow-up on seismometer discussion

The shaking started earlier today than yesterday, at ~9pm local time.

While the IFO is shaking, I thought (as Jan Harms suggested) I'd take a look at the cross-spectra between our seismometer channels at the dominant excitation frequency, which is ~1.135 Hz. Attachment #1 shows the phase of the cross spectrum taken for 10 averages (with 30mHz resolution) during the time period when the shaking was strong yesterday (~1500 seconds with 50% overlap). The logic is that we can use the relative phasing between the seismometer channels to estimate the direction of arrival and hence, the source location. However, I already see some inconsistencies - for example, the relative phase between BS_Z and EX_Z suggests that the signal arrives at the EX seismometer first. But the phasing between EX_Y and BS_Y suggests the opposite. So maybe my thinking about the problem as 3 co-located sensors measuring plane-wave disturbances originating from the same place is too simplistic? Moreover, Koji points out that for two sensors separated by ~40m, for a ground wave velocity of 1.5 km/s, the maximum phase delay we should see between sensors is 30 msec, which corresponds to ~10 degrees of phase. I guess we have to undo the effects of the phasing in the electronics chain.

Does anyone have some code that's already attempted something similar that I can put the data through? I'd like to not get sucked into writing fresh code.

🤞 this means that the shaking is over for today and I get a few hours of locking time later today evening.

Another observarion is that even after the main 1.14 Hz peak dies out, there is elevated seismic acitivity reported by the 1-3 Hz BLRMS band. This unfortunately coincides with some stack resonance, and so the arm cavity transmission reports greater RIN even after the main peak dies out. Today, it seems that all the BLRMS return to their "nominal" nighttime levels ~10 mins after the main 1.14 Hz peak dies out.

Attachment 1: seisxSpec.pdf
seisxSpec.pdf
  15035   Tue Nov 19 15:08:48 2019 gautamUpdateCDSVertex models rebooted

Jon and I were surveying the CDS situation so that he can prepare a report for discussion with Rolf/Rich about our upcoming BHD upgrade. In our poking around, we must have bumped something somewhere because the c1ioo machine went offline, and consequently, took all the vertex models out. I rebooted everything with the reboot script, everything seems to have come back smoothly. I took this opportunity to install some saturation counters for the arm servos, as we have for the CARM/DARM loops, because I want to use these for a watch script that catches when the ALS loses lock and shuts stuff off before kicking optics around needlessly. See Attachment #1 for my changes.

Attachment 1: armSat.png
armSat.png
  15034   Mon Nov 18 21:04:38 2019 gautamUpdateLSCLocking - some ideas

Some ideas Koji suggested:

  1. Try approaching the CARM offset zero point "from the other side" - i.e. start with a CARM offset of the opposite sign (I typically use negative CARM offset).
  2. With the PRMI locked, try bringing one arm onto resonance while the other arm is held off resonance. 

For the second idea, it is convenient to be able to control the arms in the XARM/YARM basis as opposed to the CARM/DARM basis as we usually do when going through the locking sequence. After some fiddling, I am able to reliably execute this transition, and achieve a state where the FP arm cavities are resonant for the IR with the ALS beat note frequency being the error signal being used. Some important differences:

  1. The frequency actuator (ETM) is weaker in this case than in the CARM/DARM basis (where MC2 is the frequency actuator) due to the longer length of the arm cavity (and for ETMX, the higher coil driver series resistance). I had to twiddle the limits of the servo banks to accommodate this. 
  2. The ALS error signal is significantly noisier than POX/POY. Hence, the control signal RMS is often in danger of saturating the DAC range. I implemented a partial fix by adding a 1st order Butterworth LPF with 1kHz corner frequency. According to the model, this eats <5 degrees of phase at the desired UGF of ~150 Hz.
  15033   Mon Nov 18 16:32:15 2019 gautamUpdateComputersZITA: started upgrade from Ubuntu 14 LTS to 18 LTS

the upgrade seems to have been successfully executed - the machine was restarted at ~430pm local time. Projector remains off and diagnostic striptools are on the samsung.

Quote:

and so it begins...until this is finished I have turned off the projector and moved the striptools to the big TV (time to look for Black Friday deals to replace the projector with a 120 inch LED TV)

  15032   Mon Nov 18 14:32:53 2019 gautamUpdatePEMFollow-up on seismometer discussion

The nightly seismic activity enhancement continued during the weekend. It always shows up around 10pm local time, persists for ~1 hour, and then goes away. This isn't a show stopper as long as it stops at some point, but it is annoying that it is eating up >1 hour of possible locking time. I walked over to CES, no one there admitted to anything - there is an "Earth Surface Dynamics Laboratory" there that runs some heavy equipment right next to us, but they claim they aren't running anything post ~530pm. Rick (building manager ?) also doesn't know of anything that turns on with the periodicity we see. He suggested contacting Watson but I have no idea who to talk to there who has an overview of what goes on in the building. 😢 

  15031   Fri Nov 15 18:59:08 2019 ranaUpdateComputersZITA: started upgrade from Ubuntu 14 LTS to 18 LTS

and so it begins...until this is finished I have turned off the projector and moved the striptools to the big TV (time to look for Black Friday deals to replace the projector with a 120 inch LED TV)

  15030   Fri Nov 15 12:16:48 2019 gautamUpdatePEMFollow-up on seismometer discussion

Attachment #1 is a spectrogram of the BS sesimometer signals for a ~24 hour period (from Wednesday night to Thursday night local time, zipped because its a large file). I've marked the nearly pure tones that show up for some time and then turn off. We need to get to the bottom of this and ideally stop it from happening at night because it is eating ~1 hour of lockable time.

We considered if we could look at the phasing between the vertex and end seismometers to localize the source of the disturbance.

Attachment 1: BS_ZspecGram.pdf.zip
  15029   Fri Nov 15 12:08:04 2019 gautamUpdateLSCPOPDC whitening board

The DC port of the Bias-Tee is routed to (a modified version of) the iLIGO whitening board. This has the well-known problem of the protection diodes of the LT1125 quad-op-amp lowering the (ideally infinite) input impedance of the first gain stage (+24 dB). To be sure as to how much signal we can put into this port (in anticipation of trying some variable finesse PRFPMI locking but also for general book-keeping), I tested the usable input range by driving a triangle wave at ~3 Hz and changing the amplitude of the signal until we observed saturation. We found that we could drive a 10 Vpp signal at which point there was evidence of some clipping (it was asymmetric, the top end of the signal was getting clipped at +14,000 cts while the bottom end still looked like a triangle wave at -16,000 counts). Anyway we probably don't want to exceed +/- 10,000 counts on this channel. This is consistent with Hartmut's statement of having +/- 4V of usable range (although the counts he mentions are twice what I saw yesterday).

Other discussion points between Rana, Koji and Gautam:

  1. Conside putting an in-vacuum (Silicon ?) QPD for the PRC angular motion sensing
    • In-vacuum will yield lower acoustic noise coupling
    • Bring the photocurrent out and do the transimpedance amplification in air 
    • Use a large area QPD so as to be more tolerant to alignment drifts without having to introduce picomotors (but how much does the POP spot actually drift and is this feasible?)
  2. Is there some better telescope configuration for the existing in-air QPD?
    • What is the correct Gouy-phase for this to be able to best sense the PRC cavity axis motion?
  15028   Fri Nov 15 11:58:12 2019 gautamUpdateLSCoff the bad Violin filters

The clue was that the loop X arm POX loop looked to have low (<3dB)) gain margin around 600 Hz (and again at 700 Hz). Attachment #1 shows a comparison of the OLTF for this loop (measured using the IN1/IN2 method) before and after our change. We hypothesize that one of the violin filters that were turned off had non-unity DC gain, because I had to lower the loop gain by 20% after these turn-offs to have the same UGF. I updated the snap files called by the arm locking scripts.

I think I caught all the places where the FM settings are saved, but some locking scripts may still try and turn on some of these filters, so let's keep an eye on it.

Quote:

We turned off many excessive violin mode bandstop filters in the LSC.

Attachment 1: violinFix.pdf
violinFix.pdf
Attachment 2: newViolinConfig.png
newViolinConfig.png
  15027   Fri Nov 15 00:18:41 2019 ranaUpdatePEMFollow-up on seismometer discussion

The large ground motion at 1 Hz started up again tonight at around 23:30. I walked around the lab and nearby buildings with a flashlight and couldn't find anything whumping. The noise is very sinusoidal and seems like it must be a 1 Hz motor rather than any natural disturbance or traffic, etc. Suspect that it is a pump in the nearby CES building which is waking up and running to fill up some liquid level. Will check out in the morning.

Estimate of displacement noise based on the observed MC_F channel showing a 25 MHz peak-peak excursion for the laser:

dL = 25e6 * (13 m / (c / lambda)

      = 1 micron

So this is a lot. Probably our pendulum is amplifying the ground motion by 10x, so I suspect a ground noise of ~0.1 micron peak-peak.

(this is a native PDF export using qtgrace rather than XMgrace. uninstall xmgrace and symlink to qtgrace.)

Attachment 1: MCshake.pdf
MCshake.pdf
  15026   Thu Nov 14 23:56:18 2019 ranaUpdateLSCoff the bad Violin filters

We turned off many excessive violin mode bandstop filters in the LSC.

Due to some feedforward work by Jenne or EQ some years ago, we have had ~10 violin notches on in the LSC between the output matrix and the outputs to the SUS.

They were eating phase, computation time, and making ~3 dB gain peaking in places where we can't afford it. I have turned them off and Gautam SDF safed it.

Offensive BS shown in brown and cooler BS shown in blue.

To rotate the DTT landscape plot to not be sideways, use this command (note that the string is 1east, not least):
pdftk in.pdf cat 1east output out.pdf

Attachment 1: out.pdf
out.pdf
  15025   Thu Nov 14 12:11:04 2019 ranaUpdatePEMFollow-up on seismometer discussion

at 1 Hz' this effect is not large so that's real translation. at lower frequencies a ground tilt couples to the horizontal sensors at first order and so the apparent signal is amplified by the double integral. drawing a free body diagram u can c that

x_apparent = (g / s^2) * theta

but for vortical this not tru because it already measures the full free fall and the tilt only shows up at 2nd order

  15024   Wed Nov 13 23:40:15 2019 gautamUpdatePEMFollow-up on seismometer discussion

Here is some disturbance in the spacetime curvature, where the local gradient of the metric seems to have been modulated (in the "downward" as well as in the other two orthogonal Cartesian directions) at ~1 Hz - seems real as far as I can tell, all the suspensions were being shaken about and all the seismometers witnessed it, though the peak is pretty narrow. A broader, less prominent peak also shows up around 0.5 Hz. We couldn't identify any clear source (no LN2 fill-up / obvious CES activity). This event lasted for ~45 mins, and stopped around 2315 local time. Shortly (~5min) after the ~1 Hz peak died down, however, the 3-10 Hz BLRMS channel reports an increase by ~factor of 2. 

Onto trying some locking now that the suspensions have settled down somewhat.

Quote:

this is due to the Equivalence Principle: local accelerations are indistinguishable from spacetime curvature. On a spherical Earth, the local gradient of the metric points in the direction towards the center of the Earth, which is colloquially known as "down".

Attachment 1: seisAll_20191111_1Hz.pdf
seisAll_20191111_1Hz.pdf
  15023   Wed Nov 13 20:15:56 2019 ranaUpdatePEMFollow-up on seismometer discussion

this is due to the Equivalence Principle: local accelerations are indistinguishable from spacetime curvature. On a spherical Earth, the local gradient of the metric points in the direction towards the center of the Earth, which is colloquially known as "down".

Quote:

I don't understand why the z-axis motion reported by the T240 is ~10x lower at 10 mHz compared to the X and Y motions. Is this some electronics noise artefact?

 

  15022   Wed Nov 13 19:34:45 2019 gautamUpdatePEMFollow-up on seismometer discussion

Attachment #1 shows the spectra of our three available seismometers over a period of ~10ksec.

  • I don't understand why the z-axis motion reported by the T240 is ~10x lower at 10 mHz compared to the X and Y motions. Is this some electronics noise artefact?
  • The difference in the low frequency (<100mHz) shapes of the T240 compared to the Guralps is presumably due to the difference in the internal preamps / readout boxes (?). I haven't checked yet.
  • There is almost certainly some issue with the EX Guralp. IIRC this is the one that had cabling issues in the past, and also is the one that was being futzed around for Tctrl, but also could be that its masses need re-centering, since it is EX_X that is showing the anomalous behaviour.
  • The coherence structure between the other pairs of sensors is consistent.

Attachment #2 shows the result of applying frequency domain Wiener filter subtraction to the POP QPD (target) with the vertex seismometer signals as witness channels.

  • The dataset was PRMI locked with the carrier resonant, ETMs misaligned.
  • The dashed lines in these plots correspond to the RMS for the solid line with the same color.
  • For both PIT and YAW, I am using BS_X and BS_Y seismometer channels for the MISO filter inputs.
  • In particular for PIT, I notice that I am unable to get the same level of performance as in the past, particularly around ~2-3 Hz.
  • The BS seismometer health indicators don't signal any obvious problems with the seismometer itself - so something has changed w.r.t. how the ground motion propagates to the PR2/PR3? Or has the seismometer sensing truly degraded? I don't think the dataset I collected was particularly bad compared to the past, and I confirmed similar performance with a separate PRMI lock from a different time period.
Attachment 1: seisAll_20191111.pdf
seisAll_20191111.pdf
Attachment 2: ffPotential.pdf
ffPotential.pdf
  15021   Thu Nov 7 17:55:37 2019 shrutiUpdateComputer Scripts / ProgramsPython packages on donatella

Today I realized that pip and other python2,3 packages were installed in the conda base environment, so after running

conda activate

I could run the python-GPIB scripts to interface with the Agilent.

Although, I did have to add a python2 kernel to jupyter/ipython, which I did in a separate conda environment:

conda create -n ipykernel_py2 python=2 ipykernel
source activate ipykernel_py2
python -m ipykernel install --user
Quote:

I've installed pyepics on Donatella running

sudo yum install pyepics

Pip and ipython did not seem to be installed yet.

 

  15020   Thu Nov 7 17:46:10 2019 shrutiUpdateALSAM measurement at X end

Some details:

- There was a SR560+SR785 (not connected for measurement) placed near the X end which I moved; it is now behind the electronics rack by the X arm beam tube (~15m away).

- Also, for the AM measurement I moved the AG5395A from behind the PSL setup to the X end, where it now is.

- By toggling the XGREEN shutter, I noticed that the cavity was not resonant before I disconnected anything from the setup since the spot shape kept changing, but I proceeded anyway. 

- Because Rana said that it was important for me to mention: the ~5 USD blue-yellow crocs (that I now use) work fine for me.

The AM Measurement:

1. The cables were calibrated with the DC block in the A port (for a A/R measurement)

2. The cable to the PZT was disconnected from the pomona box and connected to the RF out of the NA, the PD output labelled 'GREEN_REFL' was also disconnected and connected to the B port via a DC block. 

3. The ITMX was 'misaligned'. (This allowed the reflected green PD output as seen on the oscilloscope to stabilize.)

4. The PZT is modulated in frequency and the residual amplitude modulation (as observed in the measured reflected green light) is plotted, ref. Attachment 1. The parameters for the plotted data in the attachment were:

# AG4395A Measurement - Timestamp: Nov 07 2019 - 17:04:07
#---------- Measurement Parameters ------------
# Start Frequency (Hz): 10000.0, 10000.0
# Stop Frequency (Hz): 10000000.0, 10000000.0
# Frequency Points: 801, 801
# Measurement Format: LOGM, PHAS
# Measuremed Input: AR, AR
#---------- Analyzer Settings ----------
# Number of Averages: 8
# Auto Bandwidth: On, On
# IF Bandwidth: 300.0, 300.0
# Input Attenuators (R,A,B): 0dB 10dB 20dB 
# Excitation amplitude = -10.0dBm

 

 

------------------------------------

Update (19:13 7thNov19):  When the ITMX was intentionally misaligned, Rana and I checked to see if the Oplevs were turned off and they were. But while I was casually checking the Oplevs again, they were on! 

Not sure what to do about this or what caused it. 

Quote:

[Shruti, Rana]

- At the X end, we set up the network analyzer to begin measurement of the AM transfer function by actuation of the laser PZT.

- The lid of the PDH optics setup was removed to make some checks and then replaced.

- From the PDH servo electronics setup the 'GREEN_REFL' and 'TO AUX-X LASER PZT' cables were removed for the measurement and then re-attached after.

- The signal today was too low to make a real measurement of the AM transfer function, but the GPIB scripts and interfacing was tested. 

 

Attachment 1: AMTF20191107.png
AMTF20191107.png
  15019   Wed Nov 6 20:34:28 2019 KojiUpdateIOOPower combiner loss (EOM resonant box installed)

Gautam and I were talking about some modulation and demodulation and wondered what is the power combining situation for the triple resonant EOM installed 8 years ago. And we noticed that the current setup has additional ~5dB loss associated with the 3-to-1 power combiner. (Figure a)

N-to-1 broadband power combiners have an intrinsic loss of 10 log10(N). You can think about a reciprocal process (power splitting) (Figure b). The 2W input coming to the 2-port power splitter gives us two 1W outputs. The opposite process is power combining as shown in Figure c. This case, the two identical signals are the constructively added in the combiner, but the output is not 20Vpk but 14Vpk. Considering thge linearity, when one of the port is terminated, the output is going to be a half. So we expect 27dBm output for a 30dBm input (Figure d). This fact is frequently oversight particularly when one combines the signals at multiple frequencies (Figrue e). We can avoid this kind of loss by using a frequency-dependent power combiner like a diplexer or a triplexer.

Attachment 1: power_combiner.pdf
power_combiner.pdf
  Draft   Wed Nov 6 20:34:08 2019 KojiUpdateIOOEOM resonant box installed

 

Quote:

[Mirko / Kiwamu]

 The resonant box has been installed together with a 3 dB attenuator.

The demodulation phase of the MC lock was readjusted and the MC is now happily locked.

 

(Background)

We needed more modulation depth on each modulation frequency and so for the reason we installed the resonant box to amplify the signal levels.

Since the resonant box isn't impedance matched well, the box creates some amount of the RF reflections (#5339).

In order to reduce somewhat of the RF reflection we decided to put a 3 dB attenuator in between the generation box and the resonant box.

 

(what we did)

 + attached the resonant box directly to the EOM input with a short SMA connector.

 + put stacked black plates underneath the resonant box to support the wight of the box and to relief the strain on the cable between the EOM and the box.

 + put a 3 dB attenuator just after the RF power combiner to reduce RF reflections.

 + readjusted the demodulation phase of the MC lock.

 

(Adjustment of MC demodulation phase)

 The demodulation phase was readjusted by adding more cable length in the local oscillator line.

After some iterations an additional cable length of about 30 cm was inserted to maximize the Q-phase signal.

So for the MC lock we are using the Q signal, which is the same as it had been before.

 

 Before the installation of the resonant box, the amplitude of the MC PDH signal was measured in the demodulation board's monitor pins.

The amplitude was about 500 mV in peak-peak (see the attached pictures of the I-Q projection in an oscilloscope). Then after the installation the amplitude decreased to 400 mV in peak-peak.

Therefore the amplitude of the PDH signal decreased by 20 %, which is not as bad as I expected since the previous measurement indicated 40 % reduction (#2586).

 

 

  15017   Wed Nov 6 19:26:57 2019 gautamUpdatePSLSome PSL cable admin

Koji and I taked about cleaning up some of the flaky cable situation on the PSL table a while ago. The changes were implemented and are documented in Attachment #1. Now the Pomona box between the Thorlabs HV Driver and the NPRO head is sitting on the PSL table (sandwiched between some teflon pieces I found in cabinet S4 along the south arm), and the cables between these two devices are better strain relieved. I turned off the Thorlabs HV supply while working on the PMC table. The IMC could be locked after this work. Probably won't solve the long standing FSS mysteries but probably can't hurt.

Unrelated to this work: I also removed a Bias tee that was just hanging out on top of the FSS electronics, which was used for the modeSpec project.

Attachment 1: PSLcableAdmin.jpg
PSLcableAdmin.jpg
  15016   Wed Nov 6 17:45:34 2019 gautamUpdateLSC~

Here is a comparison of the response of various DoFs in our various RFPD sensors for two different CARM offsets. Even in the case of the smaller CARM offset of ~1kHz, we are several linewidths away from the resonance. Need to do some finesse modeling to make any meaningful statement about this - why is the CARM response in REFL11 apparently smaller for the smaller CARM offset?

If you mistrust my signal processing, the GPS times for which I ran the sensing lines are:

CARM offset = ~30kHz (arm transmission <0.02) --- 1257064777+5min

CARM offset = ~1kHz (arm transmission ~5) --- 1257065566+5min

Quote:

Summary:

There seems to be stronger-than-expected coupling between CARM and the 3f sensors. 

  15015   Wed Nov 6 17:05:45 2019 gautamUpdateLSCCARM calibration

Summary:

A coarse calibration of the CARM error point (when on ALS control) is 7.040 +/- 0.030 kHz/ct. This corresponds to approximately 0.95nm/ct. I typically lose the PRMI lock when the CARM offset is ~0.2 cts, which means I am about 1kHz away from the resonance. This is >10 CARM linewidths.

Details:

The calibration was done by sweeping the CARM offset (no PRM) and identifying the arm cavity FSRs by looking for peaks in TRX / TRY. Attachment #1 shows the scan, while Attachment #2 shows a linear fit to the FSRs. In Attachment #2, the frequency axis is taken from the phase tracker servo, which was calibrated by injecting a "known" frequency with the Marconi, and there is good agreement to the expected FSR with 37.79 m long arm cavities. There is much more info in the scan (e.g. modulation depths, mode matching to the arm cavities etc) which I will extract later, but if anyone wants the data (pre-downsampled by me to have a managable filesize), it's attached as a .zip file in Attachment #3.

Attachment 1: CARMscan.pdf
CARMscan.pdf
Attachment 2: CARMcalib.pdf
CARMcalib.pdf
Attachment 3: scan.hdf5.zip
  15014   Wed Nov 6 02:08:48 2019 gautamUpdateLSCLocking updates

Summary:

There seems to be stronger-than-expected coupling between CARM and the 3f sensors. 

Details:

Full analysis tomorrow, but I collected sensing matrix measurements with lines driven in PRCL,MICH and CARM at a couple of CARM offsets. I also wanted to calibrate the CARM offset to physical units so I ran some scans of the CARM offset and collected the data so I can use the arm cavity FSR to calibrate CARM. Koji suggested using REFL165_I for PRCL and REFL165_Q for MICH control - this would allow us to see if the problem was with the 1f sideband only. While the lock could be established, we still couldn't push the arm powers above 10 without breaking the PRMI lock. While changing the CARM offset, we saw a significant shift in the DC offset level of the out-of-loop REFL33_I signal. Need to think about what this means...

  15013   Tue Nov 5 12:37:50 2019 gautamUpdatePEMT240 interface unit pulled out

I removed the Trillium T240 DAQ interface unit from 1X4 for investigation.

It was returned to the electronics rack and all the connections were re-made. Some details:

  1. The board is indeed a D1000749-v2 as Koji said it is. There is just an additional board (labelled D1001872 but for which there is no schematic on the DCC) inside the 1U box that breaks out the D37 connector of the v2 into 3 D15 connectors. I took photos.
  2. Armed with the new cable Chub got, and following the manual, I ran the re-centering routine.
    • Now all the mass-monitoring position voltages are <0.3 V DC, as the manual tells me they should be.
    • I noticed that when the seismometer is just plugged in and powered, it takes a few minutes for the mass monitoring voltages to acquire their steady state values.
    • The V indicator reported ~-2V DC, and the W indicator reported -3.9V DC.
    • While running the re-centering routine, I monitored the mass-position indicator voltages (via the backplane D15 connector) on an oscilloscope. See Attachment #1 for the time series. The data was rather noisy, I don't know why this is, so I plot the raw data in light colors and a filtered version in darker colors. Also, there seems to be a gain of x2 in the voltages on the backplane relative to what the T240 manual tells me I should expect, and the values reported when I query the unit via the serial port.
    • We should ideally just install another Acromag ADC in the c1susaux box and acquire these and other available diagnostic information, since the signals are available.
    • We should also probably check the mass position indicator values in a few days to see if they've drifted off again.
    • Looking at the raw time series / spectra of the BS channels, I see no obvious signatures of any change. 
    • I will run a test by locking the PRC and looking for coherence between the seismometer data and angular motion witnessed by the POP QPD, as this was what signalled my investigation in the first place.

Update 445pm: Seems to have done something good - the old feedforward filters reduce the YAW RMS motion by a factor of a few. Pitch performance is not so good, maybe the filter needs re-training, but I see coherence, see Attachment #2 for the frequency domain WF.

Attachment 1: T240_recenter.pdf
T240_recenter.pdf
Attachment 2: ffPotential.pdf
ffPotential.pdf
  15012   Tue Nov 5 11:52:27 2019 gautamUpdateLSCLocking notes

Summary:

I am still unable to achieve arm powers greater than TRX/TRY ~10 while keeping the PRMI locked. A couple of times, I was able to get TRY ~50, but TRX stayed at ~10, or even dropped a little, suggestive of a DARM offset? On the positive side, the ALS system seems to work pretty reliably, and I can keep the arms controlled by ALS for several tens of minutes.

Details:

  • Despite my POP beam path improvements, I saw the POP22 level drop as I lowered the CARM offset.
  • One strange feature last night was that with the arms held off resonance using ALS, I had to flip the sign and increase the gain by ~x2 of the REFL33_I-->PRCL loop in order to lock the PRMI. This was confirmed by locking on the 1f error signals and measuring the ratio of the response between the 1f and 3f signals while shaking PRCL using DTT swept sine.
  • At different CARM offsets, I noted that the DC offset level on the 1f photodiodes (i.e. REFL11 and AS55) were changing significantly.
  • I ran a measurement of the sensing matrix with the arm powers hovering around ~10, which is just before I lose the PRMI lock - managed to stay locked for >5 minutes, but the sensing matrix seems to suggest that the REFL33 demod angle needs to be rotated - maybe this is the reason why the PDH horn-to-horn voltage of REFL33 is lower now than it was last week? No idea why that should be, I was around the LSC rack but if the situation is so fragile, seems hopeless.
  • MICH sensed by REFL165_Q still seems stable, so that's good...
  • So my best hypothesis at the moment is that the PRCL optical gain is falling as I reduce the CARM offset (due to DC offset? or something else?). Needs some detailed modeling for more insight, I'm out of ideas for tests to run while locking as I've gone through the full gamut of OLTF and sensing matrix measurements at various CARM offsets without getting any clues as to what's going on.
Attachment 1: PRMI3f_ALS_Nov4sensMat.pdf
PRMI3f_ALS_Nov4sensMat.pdf
  15011   Mon Nov 4 19:02:25 2019 YehonathanUpdatePSLMapping the PSL electronics

I created a spreadsheet (Attached) by taking Koji's c1psl sheet from slow_channel_list and filtering out the channels that do not need an Acromag. I added in the QPD channels that are relevant to the PSL from the c1iool0 sheet.

I began mapping the PSL related Eurocrates connectors to their respective VME channels starting with the PMC electronics.

I am confused about the TTFSS interface (D040423): While it is a Eurocrate card, in the schematics it seems to have 50 pin connectors.

I found old wiring schematics that might help with identifying the channels once the connector issue is clarified.

 

 

Attachment 1: PSL_Wirings_-_Sheet1_(1).pdf
PSL_Wirings_-_Sheet1_(1).pdf
  15010   Mon Nov 4 16:06:58 2019 gautamUpdateLSCPOP optical path

I did some re-alignment of the POP beam on the IX in air table. Here are the details:

  1. Attachment #1 - optical layout.
  2. With the PRC locked with the carrier resonant (no arm cavities), there is ~300uW of DC power incident on the Thorlabs PDA10CF, which serves as POP22, POP110 and POPDC photosensor.
    • See this elog for the signal paths.
    • On a scope, this corresponded to ~1.8 V DC of voltage. This is in good agreement with the expected transimpedance gain of 10 kOhms and responsivity of ~0.65 A/W given on the datasheet.
    • This is also in agreement with the ~6000 ADC counts I see in the CDS system (although there are large fluctuations). 
  3. These was significant misalignment of the beam on this photodiode at some point:
    • Previously, I had used the CDS system to walk the beam on thde photodiode to try and maximize the power.
    • Today I took a different approach - triggered the MICH and PRCL loops on REFLDC (instead of the usual POPDC / POP22) so I could freely block the beam.
    • I found that there is a fast (f=35mm) lens to make the beam small enough for the PDA10CF. The beam was somewhat mis-centered on this strongly curved optic, and I suspect it was amplifying small misalignments. Anyway it is much better centered now (see Attachment #2) and I have a much stronger POPDC signal (by a factor of ~2-3, see Attachment #3).
    • The ASS dither alignment now shows much more consistent behavior - minimizing REFLDC maximises POPDC, see Attachment #4.
    • I took this opportunity to take some spectra/time-series of the PD output with the interferometer in this configuration. 

Tangentially related to this work - I took the nuclear option and did a hard reboot of the c1susaux Acromag crate on Sunday to fix the EPICS issue - it seems to be gone for now, see Attachment #5.

Attachment 1: IMG_8027.JPG
IMG_8027.JPG
Attachment 2: lensRealignment.jpg
lensRealignment.jpg
Attachment 3: POPrealigned.png
POPrealigned.png
Attachment 4: POPdither.png
POPdither.png
Attachment 5: PRMfixed.png
PRMfixed.png
  15009   Mon Nov 4 15:29:47 2019 gautamUpdateLSCPOP signal path

There are many versions of the POP22 signal path I found on the elog, e.g. this thread. But what I saw at the LSC rack was not quite in agreement with any of those. So here is the latest greatest version.

Since the 2f signals are mainly indicators of power buildups and are used for triggering various PDH loops, I don't know how critical some of these things are, but here are some remarks:

  1. There is no Tee + 50 ohm terminator after the minicircuits filters, whose impedance in the stopband are High-Z (I have been told but never personally verified).
  2. The RF amplifier used is a Minicircuits ZFL-1000-LN+. This has a gain of 20dB and 1dB compression output power spec of 3dBm. So to be safe, we want to have not more than -20dBm of signal at the input. On a 50-ohm scope (AC coupled), I saw a signal that has ~100mVpp amplitude (there is a mixture of many frequencies so this is not the Vpp of a pure sinusoid). This corresponds to -16dBm. Might be cutting it a bit close even after accounting for cable loss and insertion loss of the bias tee.
  3. We use a resistive power splitter to divide the power between the POP22 and POP110 paths, which automatically throws away 50% of the RF power. A better option is the ZAPD-2-252-S+.
  4. The Thorlabs PDA10CF photodiode (not this particular one) has been modelled to have a response that can be approximated by a complex pole pair with Q=1 at ~130 MHz. But we are also using this PD for measuring the 110 MHz PD which is a bit close to the band edge?
Attachment 1: POPchain.pdf
POPchain.pdf
  15008   Mon Nov 4 13:26:04 2019 YehonathanUpdatePSLUp to date sketch of the 1x1 and 1x2 Eurocrates

Done.

Quote:

Thanks. Please update this wiki page too.

https://wiki-40m.ligo.caltech.edu/Electronics/ElectronicsRacks#A1X1

  15007   Mon Nov 4 11:41:28 2019 shrutiUpdateComputer Scripts / ProgramsEpics installed on donatella

I've installed pyepics on Donatella running

sudo yum install pyepics

Pip and ipython did not seem to be installed yet.

  15006   Sat Nov 2 17:08:34 2019 YehonathanUpdatePSLUp to date sketch of the 1x1 and 1x2 Eurocrates

Thanks. Please update this wiki page too.

https://wiki-40m.ligo.caltech.edu/Electronics/ElectronicsRacks#A1X1

  15005   Sat Nov 2 16:36:55 2019 YehonathanUpdatePSLUp to date sketch of the 1x1 and 1x2 Eurocrates

I reproduced Gautam's sketch of the 1x1 and 1x2 Eurocrates into a pdf image that contains links to the appropriate DCCs in the legend (see attachement).

Attachment 1: 1x1_1X2_Eurocrates_with_links.pdf
1x1_1X2_Eurocrates_with_links.pdf
  15004   Thu Oct 31 10:44:40 2019 gautamUpdatePSLPMC re-locked

PMC got unlocked at ~4am. I re-locked it. Also tweaked the input pointing into the cavity. The misalignment was mostly in pitch.

There was also a loud buzzing in the control room due to the audio cable being improperly seated in the mixer. I re-seated it.

  15003   Wed Oct 30 23:12:27 2019 KojiUpdateSUSPRM suspension issues

Sigh... hard loch

  15002   Wed Oct 30 19:20:27 2019 gautamUpdateSUSPRM suspension issues

While I was trying to lock the PRMI this evening, I noticed that I couldn't move the REFL beamspot on the CCD field of view by adjusting the slow bias voltages to the PRM. Other suspensions controlled by c1susaux seem to respond okay so at first glance it isn't a problem with the Acromag. Looking at the OSEM sensor input levels, I noticed that UL is much lower than the others - see Attachment #1, seems to have happened ~100 days ago. I plugged the tester box in to check if the problem is with the electronics or if this is an actual shorting of some pins on the physical OSEM as we had in the past. So PRM watchdog is shutdown for now and there is no control of the optic available as the cables are detached. I will replace the connections later in the evening.

Update 10pm:

  1. Measured coil inductances with breakout board and LCR meter - all 5 coils returned ~3.28-3.32 mH.
  2. Measured coil resistances with breakout board and DMM - all 5 coils returned ~16-17 ohms.
  3. Checked OSEM PD capacitance (with no bias voltage) using the LCR meter - each PD returned ~1nF.
  4. Checked resistance between LED Cathode and Anode for all 5 LEDs using DMM - each returned Hi-Z.
  5. Checked resistance between PD Cathode and Anode for all 5 PDs using DMM - each returned ~430 kohms.
  6. Checked that I could change the slow bias voltages and see a response at the expected pins (with the suspension disconnected).

Since I couldn't find anything wrong, I plugged the suspension back in - and voila, the suspect UL PD voltage level came back to a level consistent with the others! See Attachment #2.

Anyway, I had some hours of data with the tester box plugged in - see Attachment #3 for a comparison of the shadow sensor readout with the tester box (all black traces) vs with the suspension plugged in, local damping loops active (coloured traces). The sensing noise re-injection will depend on the specifics of the  local damping loop shapes but I suspect it will limit feedforward subtraction possibilities at low frequencies.

However, I continue to have problems aligning the optic using the slow bias sliders (but the fast ones work just fine) - problem seems to be EPICS related. In Attachment #4, I show that even though I change the soft PITCH bias voltage adjust channel for the PRM, the linked channels which control the actual voltages to the coils take several seconds to show any response, and do so asynchronously. I tried restarting the modbus process on c1susaux, but the problem persists. Perhaps it needs a reboot of the computer and/or the acromag chassis? I note that the same problem exists for the BS and PRM suspensions, but not for ITMX or ITMY (didn't check the IMC optics). Perhaps a particular Acromag DAC unit is faulty / has issues with the internal subnet?

Attachment 1: PRMUL.pdf
PRMUL.pdf
Attachment 2: PRMnormal.pdf
PRMnormal.pdf
Attachment 3: PRM-Sensors_noise.pdf
PRM-Sensors_noise.pdf PRM-Sensors_noise.pdf
Attachment 4: PRMsuspensionWonky.png
PRMsuspensionWonky.png
  15001   Wed Oct 30 17:08:40 2019 gautamUpdateLSCPOP22 investigation

The POP beam coming out of the vacuum chamber is split by a 50/50 BS and half is diverted to the POP22/POP110/POPDC photodiode (Thorlabs PDA10CF) and the other half goes to the POP QPD. This optical layout is still pretty accurate. I looked at the data of the POPDC and POP QPD SUM channels while the dither alignment was running, to see if I could figure out what's up with the weird correlated dip in REFLDC and POPDC. While the POPDC channel shows some degradation as the REFLDC level goes down (=alignment gets better), the QPD sum channel shows the expected light level increase. So it could yet be some weird clipping somewhere in the beampath - perhaps at the 50/50 BS? I will lock the PRMI (no arms) and check...

Attachment 1: POP22anomaly.pdf
POP22anomaly.pdf
  15000   Wed Oct 30 11:53:41 2019 gautamUpdateLSCMICH loop shape tuning

I changed the shape of the low pass filter to reduce high frequency sensor noise injection into the MICH control signal. The loop stability isn't adversely affected (lost ~5 degrees of phase margin but still have ~50 degrees), while the control signal RMS is reduced by ~x10. This test was done with the PRMI locked on the carrier, need to confirm that this works with the arms controlled on ALS and PRMI lcoked on sideband.

Attachment 1: MICH_ELP.pdf
MICH_ELP.pdf
Attachment 2: MICH_ELP_TFs.pdf
MICH_ELP_TFs.pdf
  14999   Wed Oct 30 01:27:00 2019 gautamUpdateLSCMore locking updates

Tried a bunch of things tonight.

  1. Modified the "ELP300" filter module in the MICH filter bank - this was really a 4th order elliptic low pass with corner at 80 Hz, which was much too low. I tried upping the corner to 500 Hz, and reducing the order, while I was able to enable the filter, there was clearly a gain-peaking feature visible after engaging this module, so the exercise of reducing the high frequency MICH actuation requires more careful (daytime) loop optimization.
  2. Tried adding some POPDC to the MICH/PRCL trigger once the PRMI was locked - I thought this would help if the problem was just with POP22 triggering turning off the MICH/PRCL loops, but the problem seems to persist with the mixed matrix trigger as well, once I reach a CARM offset where the arm powers exceed ~10, the PRMI loses lock.
  3. One strange feature I don't understand is that with the PRMI locked with the carrier field resonant, when running the dither alignment servo to minimize REFLDC (= more carrier coupled into the PRC), the POPDC level also goes down, but TRX and TRY go up slightly. I confirmed that the beam isn't falling off the POP22 photodiode (Thorlabs PDA10CF), but I don't understand why these two DC powers should fall simultaneously - if I couple more carrier into the PRC, shouldn't the POPDC level also increase?

One possibility is that the arm buildup is exerting some torque on the ITMs, which can also change the PRC cavity axis - as the buildup increases, the dominant component of the circulating field in the PRC comes from the leakage from the overcoupled arm cavity. We used to DC couple the ITM Oplev servos when locking the PRMI. The TRX level of 1 corresponds to ~5W of circulating power in the arm cavity, and the static radiation pressure force due to this circulating power is ~30 nN, rising up to 300nN as the TRX level hits 10. So for 1mm offset of the spot position on the ITM, we'd still only exert 300 pN m of torque. I don't see any transient in the Oplev error signals when locking the arm cavity as usual with POX/POY, but on timescales of several seconds, the Oplev error point shows ~3-5 urad of variation.

Attachment 1: POP_ASS.png
POP_ASS.png
  14998   Tue Oct 29 17:40:48 2019 gautamUpdateLSCMore locking updates

I set up a photodiode (PDA10CF) in the IFO REFL beampath and the Agilent NA is sitting on the east side of the PSL enclosure. This was meant to be just a first look, maybe the PDA10CF isn't suitable for this measurement. The measurement condition was - PRM aligned so we have a REFL beam (DC level = 8.4V measured with High-Z). Both ITMs and ETMs were macroscopically misaligned so that there isn't any cavity effects to consider. I collected noise around 11 and 55 MHz, and also a dark measurement, plots to follow. The optics were re-aligned to the nominal config but I left the NA on the east side of the PSL enclosure for now, in anticipation of us maybe wanting to tune something while minimizing a peak.

Attachment #1: Results of a coarse sweep from 5 MHz to 100 MHz. The broadband RIN level is not resolvable above the dark noise of the photodiode, but the peaks at the modulation frequencies (11 MHz, 55 MHz and 29.5 MHz) are clearly visible. Not sure what is the peak at ~44 MHz or 66 MHz. Come to think of it, why is the 29.5 MHz peak so prominent? The IMC cavity pole is ~4kHz so shouldn't the 29.5 MHz be attenuated by 80dB in transmission through the cavity?

Attachment #2: Zoomed in spectra with finer IF bandwidth around the RF modualtion frequencies. From this first measurement, it seems like the RIN/rad level is ~10^5, which I vaguely remember from discussions being the level which is best achieved in practise in the 40m in the past.

Quote:
 

Check the RAM due to the EOM? Perhaps the pointing / polarization control into the EOM got degraded.

Attachment 1: broadSweep.pdf
broadSweep.pdf
Attachment 2: zoomSweep.pdf
zoomSweep.pdf
  14997   Tue Oct 29 15:13:19 2019 gautamUpdateLSCMore locking updates

I looked at some signals for a 10 second period when the PRMI was locked with at some CARM offset, just before the PRMI lost lock, to see if there are any clues. I don't see any obvious signatures in this set of signals - if anything, the PRM is picking up some pitch offset, this is seen both at the Oplev error point and also in the POP QPD spot position. But why should this be happening as I reduce the CARM offset? The arm transmission is only ~5, so it's hard to imagine that the radiation pressure is somehow torquing the PRM. There are no angular feedback loops actuating on the PRM in this state except the local damping and Oplev loops.

The 1f signals are also changing their mean DC offset values, which may be a signature of a changing offset in the 3f MICH and PRCL error points? The MICH error signal is pretty noisy (maybe I can turn on some LPF to clean this up a bit), but I don't see any DC drift in the PRCL control signal.

Attachment 1: PRMI_lockloss.png
PRMI_lockloss.png
  14996   Tue Oct 29 01:24:45 2019 gautamUpdateLSCMore locking updates

Summary:

  1. The two arm lengths can be controlled reliably in the CARM/DARM basis using ALS error signals.
  2. With a CARM offset to keep the arm cavitites off resonance, the PRMI can be locked using 3f error signals.
  3. On attempting to reduce the CARM offset, I see a drop in the POP22 buildup in the PRC (correlated with the arm powers increasing). Not entirely clear why this is happening.

I ran some sensing measurements at various CARM offsets to check if the PRCL-->REFL33 and MICH-->REFL165 signals were being rotated out of the sensing quadrature as I lowered the CARM offset - there was no evidence of this happening. See Attachment #2. Other possibilities:

  • CARM offset dependant offsets in the MICH/PRCL error points?
  • Check the RAM due to the EOM? Perhaps the pointing / polarization control into the EOM got degraded.
  • Angular stability of the PRC is still pretty poor, getting the angular feedforward back up and running would help the duty cycle enormously.

The IMC went into some crazy state so I'm calling it for the night, need to think about what could be happening and take a closer look at more signals during the CARM offset reduction period for some clues...

Attachment 1: POP22_feature.png
POP22_feature.png
Attachment 2: PRMI3f_ALS_Oct21sensMat.pdf
PRMI3f_ALS_Oct21sensMat.pdf
  14995   Mon Oct 28 23:20:11 2019 gautamUpdateALSALS power budget

 

IR ALS power budget
Photodiode PSL VDC [V] PSL IDC [uA] AUX VDC [V] AUX IDC [uA] IRF [mA pk] PRF [dBm]
PSL+EX 3 300 2.5 250 ~600  ~3
PSL+EY 3 300 0.6 60 ~270 ~ -3

In calculating the above numbers, I assumed a DC transimpedance of 10 khhms and an RF Transimpedance of ~800 V/A.

[Elog14480]: per these calculations, with the NewFocus 1611 PDs, we cannot achieve shot noise limited sensing for any power below the rated maximum for linear operation (i.e. 1mW). Moreover, the noise figure of the RF amplifier we use to amplify the sensed beat note before driving the delay-line frequency discriminator is unlikely to be the limiting noise source in the current configuration. Rana suggested that we get two Gain Blocks. These can handle input powers up to ~10dBm while still giving us plenty of power to drive the delay line. This way, we can (i) not compromise on the sacred optical gain, (ii) be well below the 1dB compression point (i.e. avoid nonlinear noise effects) and (iii) achieve a better frequency discriminant

Temporary fix: While the gain blocks arrive, I inserted a 10dB (3dB) attenuator between the PSL+EX (PSL+EY) photodiode RF output and the ZHL-3A amplifiers. This way, we are well below the 1dB compression point of said RF amplifiers, and also below the 1dB compression point of the on-board Teledyne AP1053 amplifiers on the demodulator boards we use.

Nest steps: Rana is getting in touch with Rich Abbott to find out if there is any data available on the noise performance of the post-mixer IF amplifier stage in the 0.1 -30 Hz range, where the voltage and current noise of the AD829 OpAmps could be limiting the DFD performance. But in the meantime, the ALS noise seems good again, and there is no evidence of the sort of CARM/DARM coupling that motivated this investigation in the first place. Managed to execute several IR-->ALS transitions tonight in the PRFPMI locking efforts (next elog).

No new Teledyne AP1053s were harmed in this process - I'll send the 5 units back to Rich tomorrow.

  14994   Mon Oct 28 18:55:06 2019 ranaUpdateComputersrossa revival

back on new Rossa from Xi computing

  1. switched to using Display Port for video; this works. The DVi, HDMI, VGA ports are connected to the motherboard rather than the video card, so they are not active.
  2. runs super slow w/ SL 7.6; maybe some service is running after startup?
  3. install repos and update according to LLO CDS wiki
  4. add controls user and group according to LLO wiki
  5. remove gstreamer ugly because it breaks yum update
  6. run 'yum update --skip-broken' because GDS doesn't work
  7. turn off old selinux stuff
  8. modify fstab to get NFS

Next:

  1. finish mounting
  2. xfce
  3. figure out why the LLO install instructions can't install any CDS software (e.g. root, DTT, etc)

Update: Sun Nov 3 18:08:48 2019

  1. moved the SL7 fresh install repos back into etc/yum.repos.d/. The LLO instructions has me remove them, but the LLO supplied repos are no good for standard apps. After putting these back was able to install standard apps (terminator, root, diaggui)
  2. copied over /etc/fstab lines from pianosa sothat the NFS mounts work correctly
  3. added symlinks so that the NFS dirs mount in the right dirs
  4. symlink libsasl2.so.3 -> libsasl2.so.2 and now DTT runs and can get data now and in the past
  5. install XFCE
  6. sitemap / MEDM works
  7.  Did "sudo ln -s /usr/lib64/libXm.so.4 /usr/lib64/libXm.so.3" to enable StripTool.

Update: Fri Nov 15 00:00:26 2019:

  1. random hanging of machine while doing various window moving or workspace switching
  2. turned off power management in XFCE
  3. turned off power management on monitor
  4. disabled SELINUX
  5. firewalld was already off
  6. installed most, pdftk, htop, glances, qtgrace, lesstif
  7. dataviewer now works and QTgrace is much nicer than XMGrace
  14993   Fri Oct 25 01:04:49 2019 gautamUpdateALSALS electronics chain was saturating

[Koji, gautam]

Summary:

We think we got to the bottom of this issue today. The RF signal level going into the demod board is too high. This electronics chain needs some careful gain reallocation.

Details:

I was demonstrating to Koji a strange feature I had noticed in the ALS control, whereby when applying a CARM offset to detune the arms, the two arms seemed to respond differently (based on the transmission levels). This kind of CARM-->DARM coupling seemed strange to me. Anyway, I also noticed that the EPICS indicators on the ALS MEDM screen suggested ADC saturations were going on. I had never really looked at the fast time series of the inputs to the phase tracker servos, but these showed saturating behavior on ndscope traces. I went to the LSC rack and measured these on a scope, indeed, they were ~20V pp.

The output of the BeatMouth PDs are going to a ZHL-3A amplifier - we should consider replacing these with lower gain amplifiers, e.g. the Teledyne AP1053. This is relegated to a daytime task.

Other findings tonight:

While working on the PSL table, I somehow put the IMC FSS into a bad state, reminiscent of this behavior. Seems like this is linked to some flaky connection on the PSL table. One candidate is the unstable attachment of the Pomona box between the NPRO PZT and the FSS output - we should install a short BNC cable between these to avoid the lever arm situation we have right now.

  14992   Thu Oct 24 18:37:15 2019 gautamUpdatePEMT240 checkout

Summary:

The Trillium T240 seismometer needs mass re-centering. Has anyone done this before, and do we have any hardware to do this?

Details:

I went to the Trillium interface box in 1X5. In this elog, Koji says it is D1000749-v2. But looking at the connector footprint on the back panel, it is more consistent with the v1 layout. Anyway I didn't open it to check. Main point is that none of the backplane data I/O ports are used. We are digitizing (using the fast CDS system) the front panel BNC outputs for the three axes. So of the various connectors available on the interface box, we are only using the front panel DB25, the front panel BNCs, and the rear panel power.

The cable connecting this interface box to the actual seismometer is a custom one I believe. It has a 19 pin military circular type hermetic connector on one end, and a DB25 on the other. Power is supplied to the seismometer from the interface box via this cable, so in order to run the test, I had to use a DB25 breakout board to act as a feedthrough and peek at the signals while the seismometer and interface boards were connected. I used Jenne's mapping of the DB25--> 19 pin connector (which also seems consistent with the schematic). Findings:

  1. We are supplying the Trillium with 39 V DC between the +PWR and -PWR pins, while the datasheet specifies 9V to 36V DC isolated. Probably this is okay?
  2. The analog (AGND) and digital (DGND) ground pins are shorted. Is this okay?
  3. I measured the DC voltages between the AGND pin and each of the mass position outputs.
    • These are supposed to indicate when the masses need re-centering.
    • The nominal output ranges for these are +/- 4 V single-ended.
    • I measured the following values (I don't know how the U,V,W basis is mapped onto the cartesian X,Y,Z coordinates):
      U_MP: 0.708 V
      V_MP: -2.151 V
      W_MP: -3.6 V
    • So at the very least, the mass needs centering in the W direction (the manual recommends doing the re-centering procedure when one of these indicators exceeds 3.5 V in absolute value).
  4. I also checked the DC voltages of the (X,Y,Z) outputs of the seismometer on the front panel BNCs, and also on the DB25 connector (so directly from the seismometer). These are rated to have a range of 40 Vpp differential between the pins. I measured ~0V on all the three axes - this is a bit confusing as I assumed a de-centered mass would lead to saturation in one of these outouts, but maybe we are measuring velocities and not positions?
  5. We probably should consider monitoring these signals long term to inform of such drifts, what is the spare channel situation in the c1sus acromag?
  6. Interestingly, today evening, there is no excess noise in the 0.1-0.3 Hz band in the X-axis of the seismometer even though it is past 6pm PDT now, which is usually the time when the excess begins to show up. The z-axis 0.3-1Hz BLRMS channel has flatlined though...

I am holding off on attempting any re-centering, for more experienced people to comment.

  14991   Thu Oct 24 11:58:16 2019 gautamUpdateASCPRC angular feedforward

Summary:

I'd like to revive the PRC angular feedforward system. However, it looks like the coherence between the vertex seismometer channels and the PRC angular motion witness sensor (= POP QPD) is much lower than was found in the past, and hence, the stabilization potential by implementing feedforward seems limited, especially for the Pitch DoF.

Details:

I found that the old filters don't work at all - turning on the FF just increases the angular motion, I can see both the POP and REFL spots moving around a lot more on the CRT monitors.

I first thought I'd look at the frequency-domain weiner filter subtraction to get a lower bound on how much subtraction is possible. I collected ~25 minutes of data with the PRC locked with the carrier resonant (but no arm cavities). Attachment #1 shows the result of the frequency domain subtraction (the dashed lines in the top subplot are RMS). Signal processing details:

  • Data was downloaded and downsampled to 64 Hz (from 2kHz for the POP QPD signals and from 128 Hz for the seismometer signals). The 'FIR' option of scipy decimate was used.
  • FFT time used was 16 seconds for the multi-coherence calculations

The coherence between target signal (=POP QPD) and the witness channels (=seismometer channels) are much lower now than was found in the past. What could be going on here?

Attachment 1: ffPotential.pdf
ffPotential.pdf
  14990   Wed Oct 23 18:40:58 2019 gautamUpdateCDSanother round of vertex FE reboots

I wanted to restart the c1oaf model. As usual, the first time the model was restarted, it came back online with a 0x2bad error. This isn't even listed in the diagnostics manual as one of the recognized error states (unless there is a typo and they mean 0x2bad when they say 0xbad). The fix that has worked for me is to stop and start the model again, but of course, there is some chance of taking all the vertex FEs down in the process. No permutation of mxstream and daqd process restarts have cleared this error. We need some CDS/RCG support to look into this issue and fix it, it is not reasonable to go through reboots of all the vertex FEs every time we want to make a model change.

  14989   Wed Oct 23 11:49:21 2019 gautamUpdatePEMPEM BLRMS anomaly

I looked into the seismometer situation a bit more today. Here is the story so far - I think more investigation is required:

  1. There is an abrupt change in the PEM BLRMS channels around 6pm PDT every day. This has been consistently seen for the last two weeks.
  2. The seismometer spectra look normal - see Attachment #1. The reference traces are from some months ago. There is elevated activity between 0.1-0.3 Hz, but this is seen in all the seismometers in all 3 DoFs.
  3. I looked at the minute trend of the raw seismometer outputs (before being BLRMSed) for the last 200 days and don't see any abrupt change in characteristics (the data gap is due to the issue in this thread).
  4. All the correct BLRMS filters seem to be engaged in the respective filter banks.

Attachment #2 has some spectrograms (they are rather large files). They suggest that the increase in noise in the 0.1-0.3 Hz band in the BS seismometer X channel is real - but there isn't a corresponding increase in the other two seismometers, so the problem could still be electronics related.

Quote:

Yesterday, Koji and I noticed (from the wall StripTool traces) that the vertex seismometer RMS between 0.1-0.3 Hz in the X-direction increased abruptly around 6pm PDT. This morning, when I came in, I noticed that the level had settled back to the normal level. Trending the BLRMS channels over the last 24 hours, I  see that the 0.3-1 Hz band in the Z direction shows some anomalous behaviour almost in the exact same time-band. Hard to believe that any physical noise was so well aligned to the seismometer axes, I'm inclined to think this is indicative of some electronics issues with the Trillium interface unit, which has been known to be flaky in the past.

Attachment 1: seisAll_20191021.pdf
seisAll_20191021.pdf
Attachment 2: specGrams.zip
  14988   Wed Oct 23 11:14:21 2019 gautamUpdateASCPRMI ASC with QPD signals normalized.

Attachment #1 - comparison of the POP QPD PIT and YAW output signal spectra with and without them being normalized by the SUM channel. I guess the shape is different between 30-100 Hz because we have subtracted out the correlated singal due to RIN?

This did not have the effect I desired - I was hoping that by normalizing the signals, I wouldn't need to change the gain of the ASC servo as the buildup in the PRC changed, but I found that the settings that worked well for PRMI locked with the carrier resonant (no arm cavities, see Attachment #2, buildup RIN reduced by a factor of ~4) did not work for the PRMI locked with the sideband resonant. Moreover, Koji raised the point that there will be some point in the transition from arms off resonance to on resonance where the dominant field in the PRC will change from being the circulating PRC carrier to the leaking arm carrier. So the response of the actuator (PRM) to correct for the misalignment may change sign. 

In conclusion, we decided that the best approach to improve the angular stability of the PRC will be to revive the PRC angualr feedforward, which in turn requires the characterization and repair of the apparently faulty vertex seismometer.

Attachment 1: PRMI_ASC_normalization.pdf
PRMI_ASC_normalization.pdf
Attachment 2: PRMI_ASC_Oct22.pdf
PRMI_ASC_Oct22.pdf
  14987   Wed Oct 23 11:11:01 2019 gautamUpdateALSEX uPDH electronics checkout

The closest thing I can think of is here.

Quote:

Is there a loop model of green PDH that agrees with the measurement? I'm wondering if something can be done with a compensation network to up the bandwidth or if the phase lag is more like a non-invertible kind.

ELOG V3.1.3-